OpenStudy (anonymous):

hey i have no idea how 2 answer this problem, help please! here it is: A 30 foot tall tree casts a shadow that is 20 feet long. At the same time, a building casts a shadow 50 feet long. Write a proportion to find the height of the building. HELP

OpenStudy (anonymous):

no genius here but ill try to help

OpenStudy (anonymous):

I find pictures to help a lot with this..

OpenStudy (anonymous):

The ratio of height to the shadow will be same at a time for different objects in a particular location. hence as the ratio for the tree is 3 : 2 it will be same for the building too.... Hence the height of the building will be 75 feet

OpenStudy (anonymous):

You can set these problems up like a ratio. Keep the units the same on both sides (height on top/shadow on bottom) 30/12=x/50 Now cross multiply 30*50=12x Solve for x :)

OpenStudy (anonymous):

You're given to assume that the angles made by the shadows are the same.. Therefore you have two similar triangles, and the ratio of their sides will be equal.

OpenStudy (anonymous):

And no, that's not to scale since I made the shadows too long before I read the numbers you were working with.

OpenStudy (anonymous):

thanks i got it =)